Earnings Per Share Financial statement data for the years ended December 31 for Dovetail Corporation follow:...

  1. Earnings Per Share

    Financial statement data for the years ended December 31 for Dovetail Corporation follow:

    20Y3 20Y2
    Net income $823,500 $702,000
    Preferred dividends $126,000 $126,000
    Average number of common shares outstanding 75,000 shares 60,000 shares

    a. Determine the earnings per share for 20Y3 and 20Y2. Round your answers to two decimal places.

    20Y3 $ per share
    20Y2 $ per share

    b. Does the change in the earnings per share from 20Y2 to 20Y3 indicate a favorable or an unfavorable trend?
Answer #1

ANSWER:

a) Earning per share= (Net income-Preferred dividend)/Share outstanding

20Y3 = (823500-126000)/75000 = $9.3 per share

20Y2 = (702000-126000)/60000 = $9.6 per share

b) Its unfavorable trend

Note : Please post each question individually for answer as per HOMEWORKLIB RULES

_____________________________________________

If you have any query or any Explanation please ask me in the comment box, i am here to helps you.please give me positive rating.

*****************THANK YOU**************